Q3M.1 You are setting up a pair of PA speakers on a field in paration for an outdoor event. Each speaker is 0.65 m wide, and the speakers are separated by 8.2 m. To test the speakers, your coworker plays a single tone through the speakers whose frequency is 440 Hz. You are standing 52 rm directly in front of the speakers and facing them. Roughly how far would you have to walk to your left or right to hear the sound amplitude drop almost to zero? How much farther would you have to go to hear the amplitude go back to its original strength?
Answer #1

A)

Destructive interference,

dx1 = (2n + 1) lambda L/2d ...... (1)

Lambda = v/f

dx1 = (2n + 1) vL/2df

When n = 0,

dx1 = vL/2df = (343 x 52)/(2 x 8.2 x 440) = 2.47 m (Ans)

B)

dx2 = nvL/fd

When n = 1

dx2 = vL/Fd = (343 x 52)/(440 x 8.2) = 4.94 m

dx = dx2 - dx1 = 2.47 m (Ans)

Comment in case any doubt please rate my answer....